OpenStudy (anonymous):

Calculate the area of triangle CDE with altitude EF, given C (3, −2), D (−1, 2), E (2, 3), and F (0, 1). can someone hel me

OpenStudy (anonymous):

help me please

jimthompson5910 (jim_thompson5910):

hint: find the length of CD. This is the base b of the triangle. Then find the altitude EF. This is the height h of the triangle. Once you have those 2 pieces of info, you use the formula A = (b*h)/2

jimthompson5910 (jim_thompson5910):

Let me know if that helps or not

OpenStudy (anonymous):

it helped a lil but not that much can you tell me more please

jimthompson5910 (jim_thompson5910):

To find the length of CD and EF, you use the distance formula \[\Large d = \sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\]

jimthompson5910 (jim_thompson5910):

Finding the length of CD Let C = (3,-2) = (x1,y1) so x1 = 3 y1 = -2 D = (-1,2) = (x1,y1) so x2 = -1 y2 = 2 Plug all this in to get... \[\Large d = \sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\] \[\Large d = \sqrt{(3-(-1))^2+(-2-2)^2}\] \[\Large d = \sqrt{(3+1)^2+(-2-2)^2}\] \[\Large d = ???\]

OpenStudy (anonymous):

ok i understand how to do that but theres four coordinates these are the answers 4 square units 6.2 square units 8 square units 8.7 square units im thinking b though

jimthompson5910 (jim_thompson5910):

what's the length of CD?

jimthompson5910 (jim_thompson5910):

in radical form

OpenStudy (anonymous):

32 so then you square root it?

OpenStudy (anonymous):

is that rite?

jimthompson5910 (jim_thompson5910):

good, so CD is \(\Large \sqrt{32}\) units long. This is the length of the base. What's the height?

OpenStudy (anonymous):

how do i do that now? i need to do the same thing again?

jimthompson5910 (jim_thompson5910):

yes, but now using points E and F

jimthompson5910 (jim_thompson5910):

since EF is the altitude (ie height) for base CD

OpenStudy (anonymous):

ok so now i got 20 what do i do now?

jimthompson5910 (jim_thompson5910):

that is incorrect

jimthompson5910 (jim_thompson5910):

we have E (2, 3), and F (0, 1) so make (x1,y1) = (2,3) (x2,y2) = (0,1)

jimthompson5910 (jim_thompson5910):

tell me what you get when you plug that into \[\Large d = \sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\]

OpenStudy (anonymous):

oohhh im sorry i got 8 now

OpenStudy (anonymous):

now wat do i do?

jimthompson5910 (jim_thompson5910):

should be \(\Large \sqrt{8}\)

jimthompson5910 (jim_thompson5910):

\[\Large A = \frac{b*h}{2}\] \[\Large A = \frac{(\sqrt{32})*(\sqrt{8})}{2}\] \[\Large A = ???\]

OpenStudy (anonymous):

ok so its 8?

jimthompson5910 (jim_thompson5910):

Correct, the answer is 8 square units